### v2

# Inversion-DPO: Precise and Efficient Post-Training for Diffusion Models

Abstract
Recent advancements in diffusion models (DMs) have been propelled by alignment methods that post-train models to better conform to human preferences. However, these approaches typically require computation-intensive training of a base model and a reward model, which not only incurs substantial computational overhead but may also compromise model accuracy and training efficiency. To address these limitations, we propose Inversion-DPO, a novel alignment framework that circumvents reward modeling by reformulating Direct Preference Optimization (DPO) with DDIM inversion for DMs. Our method conducts intractable posterior sampling in Diffusion-DPO with the deterministic inversion from winning and losing samples to noise and thus derive a new post-training paradigm. This paradigm eliminates the need for auxiliary reward models or inaccurate appromixation, significantly enhancing both precision and efficiency of training. We apply Inversion-DPO to a basic task of text-to-image generation and a challenging task of compositional image generation. Extensive experiments show substantial performance improvements achieved by Inversion-DPO compared to existing post-training methods and highlight the ability of the trained generative models to generate high-fidelity compositionally coherent images. For the post-training of compostitional image geneation, we curate a paired dataset consisting of 11,140 images with complex structural annotations and comprehensive scores, designed to enhance the compositional capabilities of generative models. Inversion-DPO explores a new avenue for efficient, high-precision alignment in diffusion models, advancing their applicability to complex realistic generation tasks. Our code is available at https://github.com/MIGHTYEZ/Inversion-DPO
1 Introduction
Diffusion models (DMs) [48, 46, 7, 6, 53, 76, 32, 43, 21] have achieved state-of-the-art performance in image generation. This advancement is largely attributed to recent developments in alignment methods of DMs [22, 55, 28, 8, 72, 63, 29], which enable models to be fine-tuned to better align with specific standards or human preferences. These methods typically require training two models simultaneously: a base model and a reward model. While this training paradigm incentivizes superior model performance, it also introduces significant computational overhead.
Diffusion-DPO [57] reparameterizes the reward score into the loss function of the base diffusion model, eliminating the need for an auxiliary model and establishing a new paradigm for DMs to learn from human preferences. However, we identify a post-training issue in Diffusion-DPO, primarily due to its approximation of the reverse process using the forward distribution during derivation. This approximation leads to suboptimal computational accuracy and inefficient training.
To overcome this limitation, we draw insights from DDIM inversion [52] to establish a precise alignment framework for diffusion model fine-tuning. Our Inversion-DPO method fundamentally enhances approximation accuracy by adopting DDIM Inversion to approximate the reverse process $p_{\theta}(x_{1:T}|x_{0})$. This proposal eliminates the distributional mismatch inherent in previous approaches, improving approximation accuracy of sampling trajectories reversely from a sample $x_{0}$ to inversion samples $x_{1:T}$.
Such precision further translates to substantial efficiency gains through simplification of the final loss. By leveraging the trajectory of DDIM Inversion, our method reduces the original four KL divergence terms in Diffusion-DPO’s derivation to two matching terms between aligned model predictions and reference model prediction. This enables more than 2$\times$ faster training convergence compared to baseline implementations. The efficiency improvements are particularly crucial for large-scale diffusion models like SDXL [43], where traditional approximation methods incur high computational costs.
To achieve high-quality and reliable image generation, we generalize Inversion-DPO to two specific tasks: basic text-to-image generation and compositional image generation. Extensive experiments demonstrate the effectiveness of our method in both diffusion model training and inference, as well as its superior performance compared with current SOTA methods, achieving significant improvements in key evaluation metrics (mention the specific improvements, e.g., percentage gains). LABEL:fig:_teaser presents representative generated samples. Furthermore, we introduce a novel dataset with structural annotations and complex scene preference scores, enabling structured evaluation and deeper insight into Inversion-DPO’s compositional preference learning.
In summary, by leveraging the principles of DDIM inversion, we address the computational accuracy and efficiency limitations of DPO-based methods. The optimized diffusion model achieves state-of-the-art performance in both basic text-to-image and compositional image generation tasks.
2 Related Work
Alignment of Diffusion Models. As generative models are widely used, aligning outputs with user preferences has become a key focus. Researchers are integrating Reinforcement Learning from Human Feedback (RLHF) [42, 2] into diffusion models to improve controllability and accuracy. A common approach involves using pre-trained or additional reward models [22, 55, 28, 8, 72], leveraging human feedback and external signals to fine-tune generation. ImageReward [63] trains a reward model based on human feedback to improve image quality and semantic alignment. DPOK [11] and DDPO [3] combine reward signals with reinforcement learning for fine-tuning diffusion models, guiding generation through real-time feedback. ReNO [10] introduces a strategy that adjusts initial noise with reward signals, bypassing model training. IterComp [71] collects and combines model preferences, using iterative feedback to enhance generative capability.
The methods above typically require explicit training of reward models. Alternatively, some approaches bypass reward models and directly fine-tune generative models via reinforcement learning. [40] For example, Diffusion-DPO [57] simplifies training by using differentiable reward signals for end-to-end fine-tuning. D3PO [65] eliminates explicit reward model training by using human feedback to guide the process. A Dense Reward View [68] enhances reward fine-tuning with dense feedback at each denoising step, improving efficiency and stability. SPO [33] adjusts denoising performance at each step with a preference model and resampler. While these DPO-based methods reduce reward model overhead, they still face post-training challenges. Recent works have explored using inversion techniques to improve DPO for diffusion models. DDIM-InPO [37] employs a reparameterization approach that treats diffusion models as single-step generators, selectively fine-tuning only the latent variables most correlated with preference data. SmPO-Diffusion [38]introduces smoothed preference distributions combined with a Renoise Inversion technique, performing multiple DDIM inversion steps followed by an additional renoise step to handle preference variability. In contrast, our Inversion-DPO directly leverages complete DDIM inversion to precisely compute the intractable posterior $p_{\theta}(x_{1:T}|x_{0})$, eliminating the approximation errors inherent in previous methods and achieving a more efficient loss formulation with only two matching terms instead of four KL divergences, improving accuracy and efficiency.
Compositional Image Generation. Compositional Text-to-Image Generation is a challenge, especially when involving multiple objects and complex relationships [67, 20]. While many studies have improved generative models, the sequential text format still limits results, particularly in compositional image generation. Researchers have proposed methods to enhance control and spatial awareness. For instance, Compositional Diffusion [35] and Attend-and-Excite [5] improve generation efficiency but face limitations in complex scenes. Methods like GLIGEN [30] and Ranni [13] incorporate spatial conditioning and multimodal information to improve control, though they come with high training costs. BoxDiff [62] and RealCompo [70] optimize cross-attention and balance realism and complexity but remain dependent on bounding box accuracy. MIGC [74] and MIGC++ [73] address multi-instance composition using multimodal descriptions but focus on spatial control, failing to fully address abstract semantic relationships between objects.
To overcome the limitations of text formats, some studies have focused on using scene graphs for compositional generation [23, 1, 35, 67, 60, 15]. Scene graphs (SG) consist of nodes and edges representing objects and their relationships. Compositional SG2IM methods aim to generate high-quality images with multiple objects and complex relationships by better understanding and combining scene elements [12, 58, 61].
SGDiff [66] enhances scene graph-based image generation by pretraining a scene graph encoder and integrating it with Stable Diffusion. SG-Adapter [51] fine-tunes Stable Diffusion to incorporate scene graph information, improving image quality and semantic consistency. R3CD [34] introduces SG Transformers to expand diffusion models, learning abstract object interactions in larger datasets. DisCo [59] combines scene graph decomposition with VAEs and diffusion models for more diverse outputs. The LAION-SG dataset [31] further improves the model’s understanding of complex scenes.
Despite progress, compositional image generation still faces challenges in fidelity and efficiency. Our proposed method shows superior performance on difficult downstream tasks.
3 Preliminary
3.1 Diffusion-DPO
Reinforcement Learning from Human Feedback (RLHF) typically involves two stages: training a reward model from pairwise preferences and optimizing policies via reinforcement learning (RL). Direct Preference Optimization (DPO) [45] bypasses this complexity by directly optimizing policies using preference data under a classification objective. Formally, given preference pairs $(x_{0}^{w},x_{0}^{l})$ conditioned on prompts $c$, the Bradley-Terry model defines preference likelihood as:
$p_{\text{BT}}(x_{0}^{w}\succ x_{0}^{l}|c)=\sigma\left(r(c,x_{0}^{w})-r(c,x_{0}^{l})\right),$ (1)
Here $r(\cdot,\cdot)$ is a reward function. Traditional RLHF maximizes the KL-regularized reward:
$\max_{p_{\theta}}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(c,x_{0})]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0}|c)\|p_{\theta_{0}}(x_{0}|c)],$ (2)
with $p_{\theta}$ as the distribution to be optimized and $p_{\theta_{0}}$ a reference distribution. DPO reparameterizes the optimal policy $p_{\theta}^{*}$ as:
$p_{\theta}^{*}(x_{0}|c)\propto p_{\theta_{0}}(x_{0}|c)\exp\left(r(c,x_{0})/\beta\right).$ (3)
Substituting this into the Bradley-Terry likelihood eliminates the reward function, yielding the DPO loss:
$\mathcal{L}_{\text{DPO}}=-\mathbb{E}_{(c,x_{0}^{w},x_{0}^{l})}\left[\log\sigma\left(\beta\log\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\theta_{0}}(x_{0}^{w}|c)}-\beta\log\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\theta_{0}}(x_{0}^{l}|c)}\right)\right].$ (4)
This implicitly optimizes rewards while maintaining policy proximity to $p_{\theta_{0}}$, avoiding explicit reward modeling and RL instability.
For diffusion models, the challenge lies in defining likelihoods over generated images $x_{0}$. Diffusion-DPO [57] extends DPO by leveraging the evidence lower bound (ELBO) of diffusion processes. The KL-regularized reward maximization objective becomes:
$\max_{p_{\theta}}\mathbb{E}_{p_{\theta}(x_{1:T}|c,x_{0})}\left[R(c,x_{0:T})\right]-\beta\mathbb{D}_{\text{KL}}[p_{\theta}(x_{0:T}|c)\|p_{\theta_{0}}(x_{0:T}|c)],$ (5)
Here $R(c,x_{0:T})$ denotes the reward over the full diffusion trajectory. Given a wining or losing sample $x_{0}$, a posterior sampling $p_{\theta}(x_{1:T}|x_{0},c)$ is required to gain the noisy sample sequence $x_{1:T}$ of $p_{\theta}$, which is intractable. By approximating the reverse process with the forward process $q(x_{1:T}|x_{0})$, the loss simplifies to:
$\mathcal{L}_{\text{Diffusion-DPO}}=-\mathbb{E}_{(x_{0}^{w},x_{0}^{l}),t,q(x_{t}^{w}\mid x_{0}^{w}),q(x_{t}^{l}\mid x_{0}^{l})}\log\sigma(-\beta T\omega(\lambda_{t})($
$\|\epsilon^{w}-\epsilon_{\theta}(x_{t}^{w},t)\|_{2}^{2}-\|\epsilon^{w}-\epsilon_{\theta_{0}}(x_{t}^{w},t)\|_{2}^{2}-$
$(\|\epsilon^{l}-\epsilon_{\theta}(x_{t}^{l},t)\|_{2}^{2}-\|\epsilon^{l}-\epsilon_{\theta_{0}}(x_{t}^{l},t)\|_{2}^{2}))),$ (6)
Here $\epsilon_{\theta}$ is the denoising network, $\epsilon_{\theta_{0}}$ the pretrained network, and $\omega(\lambda_{t})$ absorbs all coefficients. This formulation improves $\epsilon_{\theta}$ on denoising of preferred samples $x_{0}^{w}$ more than dispreferred $x_{0}^{l}$, aligning diffusion models with preferences without additional inference costs or mode collapse.
3.2 DDIM Inversion
In this part, we introduce DDIM Inversion which is crucial in our proposed method. In traditional diffusion models like DDPM [17], the generative process starts with Gaussian noise $x_{T}$ and gradually denoises through a reverse Markov chain to obtain a clean image $x_{0}$. Although the training procedure maximizes a variational lower bound, the sampling process is inherently stochastic, making it difficult to precisely recover the sampling path $\{x_{t}\}$ for a given $x_{0}$.
Denoising Diffusion Implicit Models (DDIM) [52] proposes a non-Markovian inference process. Given any intermediate state $x_{t}$, the trained $\epsilon_{\theta}(x_{t},t)$ estimates the corresponding denoised sample:
$\hat{x}_{0}=f_{\theta}^{(t)}(x_{t})=\frac{x_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}(x_{t},t)}{\sqrt{\alpha_{t}}},$ (7)
where $\alpha_{t}$ is a signal retention coefficient associated with time step $t$. When $t>1$, the generative process is given by
$\begin{split}p_{\theta}(x_{t-1}|x_{t})&=\mathcal{N}\left(\sqrt{\alpha_{t-1}}\hat{x}_{0}\right.\\ &\quad+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\cdot\frac{\boldsymbol{x}_{t}-\sqrt{\alpha_{t}}\hat{x}_{0}}{\sqrt{1-\alpha_{t}}},\\ &\quad\left.\sigma_{t}^{2}\boldsymbol{I}\right)\end{split}$ (8)
When $t=1$, $p_{\theta}(x_{0}|x_{1})=\mathcal{N}\left(\hat{x}_{0},\sigma_{1}^{2}\boldsymbol{I}\right)$. The noise term $\sigma_{t}$ can be zero during sampling, thus converting the generative process into a deterministic implicit probabilistic model. Since the update process no longer introduces additional randomness, DDIM not only allows high-quality samples to be generated in fewer steps but also renders the generative mapping invertible, which enables recovering the sampling trajectory $\{x_{1},x_{2},\ldots,x_{T}\}$ from $x_{0}$ [52].
It has been observed that when the number of sampling steps is large, predicted noises are close in adjacent steps [75]. This supports the assumption that the ordinary differential equation (ODE) process can be reversed within the limit of small steps [25], which means $\epsilon_{\theta}(x_{t},t)\approx\epsilon_{\theta}(x_{t-1},t-1)$. This key property allows us to invert the DDIM generative process [9, 24]. Starting from a real sample $x_{0}$, we can backtrack to sampling its potential denoising sequence $\{x_{1},x_{2},\ldots,x_{T}\}$. The detailed inversion iteration is
$x_{t}=\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}}x_{t-1}+(\sqrt{1-\alpha_{t}}-\sqrt{\frac{\alpha_{t}}{\alpha_{t-1}}-\alpha_{t}})\epsilon_{\theta}(x_{t-1},t-1)$ (9)
Such iteration finally reaches $x_{T}$. Under the assumption above, with $x_{T}$ as the initial noise, a deterministic sampling process $\epsilon_{\theta}$ is likely to recover the sequence $\{x_{T-1},\ldots,x_{1}\}$ and come back to the original $x_{0}$. More advanced estimations [14, 24] give more precise recovery. This capability to recover the latent path from $x_{0}$ provides a theoretical basis for Inversion-DPO we introduce next.
4 method
Figure 1: Computation pipeline of the proposed Inversion-DPO. 1) Given an initial image $x_{0}$, we obtain the diffusion trajectories $\{x_{1},x_{2},\ldots,x_{T}\}$ of both the winning and losing samples from a reference model via DDIM inversion. 2) At each timestep along the diffusion trajectory, the model is trained to predict the noise $\epsilon$, aligning it closely with the $\epsilon$ of the winning sample while diverging from that of the losing sample. 3) This process is repeated across the entire diffusion trajectory, continuously updating model parameters until the final predicted image is generated.
4.1 Inversion Direct Preference Optimization
DPO [45] is an effective reward-guided post-training approach that directly learns human preferences through a reparameterized policy, rather than optimizing a reward function and then performing reinforcement learning (RL).
Recall that when adapting DPO to diffusion models [57] , the objective shifts from fitting $x_{0}$ to fitting the trajectory $(x_{0},\dots,x_{T})$with
$(\displaystyle\mathcal{L}_{\text{Diffusion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}$10)
$\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}\left[\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}-\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{l})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{l})}\right]\bigg{)}$
To optimize this, $\boldsymbol{x}_{1:T}$ across all time steps are required with a posterior sampling of $p_{\theta}$. Thus, Diffusion-DPO encounters two key limitations: 1) The intractability of $p_{\theta}(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})$ results in a subtitution with $q(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})$ as approximation, introducing large estimation errors; 2) The KL-Divergence formulation expands the original pairwise comparison into four terms.
We address these issues by designing Inversion-DPO (Fig. 1). Our key insight stems from the deterministic characteristic of DDIM Inversion [52]. Given a sample $\boldsymbol{x}_{0}$, DDIM Inversion constructs the sampling trajectory $\boldsymbol{x}_{1:T}$ leading to $\boldsymbol{x}_{0}$ through a deterministic mapping by equation (9). Thus, we propose to approximate the posterior $p_{\theta}(\boldsymbol{x}_{1:T}|\boldsymbol{x}_{0})$ with DDIM inversion. This eliminates the need for stochastic approximation $q$ in trajectory estimation. Using DDIM Inversion is reasonable with three advantages. First, assuming the deterministic sampling setting, the posterior inference is the inversion process. Thus, the resulting $\boldsymbol{x}_{1:T}$ are more precise. Second, such inversion is compatible with the definition of the reward function on diffusion in equation (5), so the implicit reward learning is also precise. Third, the final estimated expectation in (10) is also more precise in this case.
Given the $\boldsymbol{x}_{1:T}^{w}$, we derive the following computation within the expectation. Specifically, $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}$ is first decomposed into the differences in conditional probabilities at each time step:
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$ (11)
Since $p_{\theta}$ and $p_{\theta_{0}}$ share the same distribution of $x_{T}$ as standard Gaussian, $\log p_{\theta}(x_{T}^{w})-\log p_{\theta_{0}}(x_{T}^{w})$ cancel out. The conditional probabilities of $p_{\theta}$ and $p_{\theta_{0}}$ are Gaussian distributions as in equation (8) and becomes dirac-delta as $\sigma_{t}$ approaches 00. We first continue derivation with $\sigma_{t}\neq 0$ and adopt a simple approximation [17] of the final loss.
Given the assumption of inversion in Sec. 3.2, the values of $\boldsymbol{x}_{t-1}^{w}$ are similar in the inversion from $x_{0}^{w}$ and in the sampling from $x_{T}^{w}$. Therefore, $\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ is only related to $\sigma_{t}^{2}$ and thus constant. At the same time, given
$\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}+$ (12)
$\left(\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}-\frac{\sqrt{\alpha_{t-1}({1-\alpha_{t}})}}{\sqrt{\alpha_{t}}}\right)\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)$ (13)
we have the following (details included in supplementary material):
$-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ $\propto\|\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})-\boldsymbol{x}_{t-1}^{w}\|^{2}$ (14)
$\propto\|\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)\|^{2}$
Summing the contributions across all time steps, we can arrive at:
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}\propto\left[\sum_{t=1}^{T}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}\right],$ (15)
The term for the losing sample in equation (10) is derived analogously. Each term in the objective can be computed exactly. The log-probability ratio now directly captures the policy divergence without introducing approximation errors.
Taking all together, the final objective becomes:
$\mathcal{L}_{\text{Inversion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}$ (16)
$\sum_{t=1}^{T}\Bigg{[}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}-\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{l},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{l},t)\|^{2}\Bigg{]}\bigg{)}$
Notably, during the training of Inversion-DPO, the previous four-term loss collapses into two deterministic summations due to trajectory determinism, which is conceptually simpler. This simplification also yields non-negligible improvements in optimization efficiency for model training. The pretrained $\epsilon_{\theta_{0}}$ can be transformed to approximate the score. Fitting the noise added as in (3.1) introduces extra variance [56, 41] which in unstable in the training [64], while fitting the predicted noise close to the score is more stable and thus improves the training efficiency. Our training schema fundamentally resolves the precision-efficiency tradeoff in preference-based diffusion alignment.
4.2 Inversion-DPO with multiple objectives
For effective Inversion-DPO training, both positive (i.e., winning samples) and negative (i.e., losing samples) data pairs are required to guide the optimization. While manual annotation is commonly used to obtain such pairs [71, 57], it is costly and prone to ambiguity. To address this, recent works [54, 3, 65, 39] adopt automatic metrics to rank images as guidance. We follow this strategy. However, relying on a single metric is often inadequate to comprehensively capture the quality of the outputs, especially in complex scene generation scenarios. Therefore, we introduce a multi-objective learning strategy, leveraging multiple metrics to provide more reliable and informative preference signals.
Specifically, we incorporate multiple evaluation metrics to derive the reward scores for data pairs, computed as follows:
$r(c,x_{0})=\frac{1}{N}\sum_{i=1}^{N}r_{i}\left(c,x_{0}\right)$ (17)
Here, $N$ denotes the number of evaluation metrics utilized in the current task, and $(c,x_{0})$ represents the conditional sample pair generated by $p_{\theta_{0}}$ given the condition $c$.
We evaluate the effectiveness of Inversion-DPO in two representative tasks. In addition to the typical text-to-image generation, we further assess it in the more challenging setting of compositional image generation, which requires the model to accurately synthesize complex scenes involving multiple objects and relationships. Existing studies [31, 59] suggest that, compared to textual annotations, structured annotations provide a more compact and explicit form of information integration, making them more effective for representing complex scenes. Such structured annotations are typically represented as graphs, consisting of nodes and edges that denote objects in the scene and their interactions, respectively.
On basic text-to-image tasks, Pick-a-Pic dataset [26] serves as a reliable source of supervision for training Inversion-DPO. However, for compositional generation, there is a lack of corresponding paired datasets that can be directly utilized. This is because existing datasets with structured annotations [4, 27, 31] provide only single-condition samples, making them insufficient for preference-based learning. To bridge this gap, we construct a paired dataset with structured annotations to support alignment and evaluation.
In detail, for a given complex scene annotation, we generate four samples using an existing baseline [31], and compute a multi-dimensional score for each generated image. During training, the model determines positive and negative samples based on a combination of these scores, guiding the learning of Inversion-DPO.
To ensure effective evaluation of complex scene images, we adopt two representative metrics–DPG-Bench Score [18] and the 3-in-1 score from T2I-CompBench [19]–as the criteria for multi-objective learning. In total, we generate 11,140 images for 2,785 annotations, with each image associated with a comprehensive complex scene score. Rather than enforcing fixed pairwise bindings, we allow the model to dynamically determine relative “winning” and “losing” samples based on their comprehensive scores during training. This approach prevents the dataset size from being constrained to the original 2,785 samples, enabling scalable data expansion within a certain range and ultimately improving training quality.
5 Experiments
Figure 2: A demonstration of how PickScore increase with more training steps. The blue curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our retraining with the official implementation of Diffusion-DPO [57], while the dashed line corresponds to the officially released weights. Our model with 80-step inversion outperforms the both the retraining and the released Diffusion-DPO checkpoint. All models are trained under identical configurations.
5.1 Dataset and Implementation Details
Dataset. We use the Pick-a-Pic v2 dataset [26] to train our T2I model, which contains paired preferences of images generated by SDXL-beta and Dreamlike, along with prompts and preference pairs collected from the web. Following the same data processing approach as Diffusion-DPO [57], we exclud 12% of tie samples, ultimately obtaining 851,293 effective data pairs. For the compositional image generation task, as described in Section 4.2, we curate compositional image pairs from LAION-SG [31] for training, ensuring data quality and diversity.
Training. During the training process, we employ the Adafactor [50] optimizer for SDXL and SGXL-SG to save memory, while AdamW [36] was used for SD1.5 training. The learning rate for all models was uniformly set to 1e-6. Specifically, the $\beta$ parameter in Inversion-DPO was set to 2000. We also set the Inversion Steps to 20, 40, and 80, respectively, to observe the impact of different DDIM Inversion steps on model performance, which is detailed in our ablation study. Each model was trained for 2000 steps, with all training tasks completed on two NVIDIA H800 GPUs.
5.2 Baselines and Evaluation Metrics
For the typical text-to-image generation task, we evaluate a set of representative baseline models and post-training methods, including SD1.5 [47], SDXL [43], Diffusion-DPO [57], DDPO [3], D3PO [65], Demon [69], and IterComp [71]. The generated results are assessed using PickScore [26], CLIP Score [44], and Aesthetic Score [49], which respectively measure alignment with human preference, semantic consistency with the input prompt, and the aesthetic quality of image details.
For compositional image generation, we evaluate SGDiff [66], SG-Adapter [51], SDXL-SG [31], and SD1.5-SG [31]. Considering the objectives of complex scene generation, we adopt FID [16] as the image quality metric, along with SG-IoU, Entity-IoU, and Relation-IoU [51], which respectively assess scene-level consistency, object-level consistency, and the consistency of relationships between objects within the generated images.
5.3 Qualitative Results
Figure 3: A comparison of generation details. Existing models often struggle to produce fine-grained details, while Inversion-DPO demonstrates a higher level of fidelity in generating nuanced elements such as limbs, ties, and textures.
Figure 4: Qualitative results on compositional image generation. Inversion-DPO proves effective in addressing key challenges in compositional image generation, including object omission (orange and green annotations in (a), (d), and (e)), attribute errors (red annotations in (b), (c), and (d)), and relation inaccuracies (blue annotations in (c) and (d)). As shown in the last row, Inversion-DPO achieves satisfactory generation results on cases where baseline models struggle.
LABEL:fig:_teaser illustrates the visual performance of our Inversion-DPO compared to baseline models on text-to-image tasks. Each row shows an input prompt and the corresponding generated images. Compared to existing baselines [43, 3, 69, 57], Inversion-DPO exhibits higher visual appeal and achieves higher PickScore. Moreover, in Fig. 3, our method demonstrates promising potential in controlling fine-grained detail generation, which is a persistent challenge for text-to-image models. We further visualize results on more challenging compositional image generation tasks. As shown in Fig. 4, Inversion-DPO demonstrates superior performance in terms of object occurrence, relation accuracy, and attribute binding.
We also conduct a user study, and Inversion-DPO enjoys 66.53% human preference compared to SDXL and 72.67% versus Diffusion-DPO. More details are in supplementary material.
5.4 Quantitative Results
Method PickScore CLIP Aesthetic
SDXL [43] 0.223 0.334 6.07
Diffusion-DPO [57] 0.227 0.340 6.05
DDPO [3] 0.222 0.336 6.10
D3PO [65] 0.223 0.338 5.99
Demon [69] 0.220 0.336 6.15
IterComp [71] 0.232 0.340 6.22
Inversion-DPO 0.232 0.341 6.24
SD1.5 [47] 0.207 0.320 5.61
[…]
Table 1: Quantitative results of typical text-to-image generation. $\mathcal{y}$ indicates that the model is built upon SD1.5. The same applies to the following tables.
Method FID SG-IoU Ent.-IoU Rel.-IoU
SGDiff [66] 35.8 0.304 0.787 0.698
SG-Adapter [51] 27.8 0.314 0.771 0.693
R3CD [34] 27.0 0.342 0.803 0.715
SDXL-SG [31] 26.7 0.340 0.792 0.703
[…]
SD1.5-SG 56.3 0.179 0.614 0.530
[…]
Table 2: Quantitative results on the more challenging task of compositional image generation. $\ast$ means the results of Inversion-DPO trained on the compositional paired dataset.
Table 1 reports the quantitative performance on typical text-to-image generation for both baseline models and ours. Compared to base T2I models [43, 47], post-training methods [57, 3, 65, 71], and training-free approach [69], Inversion-DPO achieves state-of-the-art results in terms of human preference alignment (PickScore), text-image consistency (CLIP Score), and aesthetic quality (Aesthetic Score). In addition to the formal SDXL-based version of Inversion-DPO, we also validate its adaptability on SD1.5.
We further investigate the performance of the proposed method on compositional image generation. As in Tab. 2, after fine-tuning on the compositional dataset, Inversion-DPO achieves the best image quality as measured by FID, outperforming existing baseline models. Moreover, it obtains the highest scores on complexity evaluation metrics, which comprehensively assess the fidelity of the generated images at the scene, entity, and relation levels.
Empirically, our method also brings notable improvements in training efficiency. In Fig. 2, Inversion-DPO achieves superior performance with significantly fewer training steps compared to Diffusion-DPO. Due to hardware limitation, our reimplementation of Diffusion-DPO with the official implementation does match the performance of the official checkpoint. Besides, we observe a slight decay of performance at the begining which may be a result of the instability of Diffusion-DPO’s training. Nevertheless, under the same resources, our Inversion-DPO constantly improves the performance of the generative model. Specifically, our variants with only 20-step DDIM Inversion reaches a performance of 0.2247 at 1000 steps, comparable to Diffusion-DPO’s 0.2245 at 2000, over 2$\times$ faster. Inversion-DPO with 80 steps is over 4$\times$ faster. Furthermore, at 2000 steps Inversion-DPO with 80-step inversion also outperforms the reported performance of Diffusion-DPO as labeled by the dotted line, further demonstrating the advantages of our method in both precision and efficiency. Results on other metrics are in Supplementary Material.
5.5 Ablation Study
# Inversion Steps PickScore CLIP Aesthetic
20 0.228 0.340 6.08
40 0.229 0.341 6.13
80 0.232 0.341 6.24
Table 3: Effect of inversion steps on model performance.
Analysis of the different inversion steps. We analyze the impact of different DDIM inversion steps on model training, with results shown in Tab. 3 and Fig. 2. It is evident that as the number of inversion steps increases, the model consistently achieves better performance across all metrics. This can be attributed to the fact that with more DDIM inversion steps, the distributional gap between adjacent steps becomes smaller, satisfying the assumption in Sec. 3.2 and making it easier to guide the model toward correct optimization directions. The configuration with 80 inversion steps is ultimately selected for training Inversion-DPO for a balance between performance and computational cost.
6 Conclusion
We presents Inversion-DPO, a novel post-training framework that integrates DDIM Inversion with DPO to achieve efficient and precise alignment in diffusion models. By leveraging the deterministic trajectory recovery of DDIM inversion, our method reduces approximation errors, achieving more than 2$\times$ training acceleration. A multi-objective scoring strategy further enhances robustness in compositional image generation. Inversion-DPO achieves SOTA performance on both text-to-image generation and compositional image generation. We additionally curate a structured dataset of 11,140 annotated images to support complex scene synthesis.
Supplementary Material
S7 User Study
To evaluate the alignment between the generated image and human cognition, we design a user study involving two control groups: the group of SDXL [43] vs. Inversion DPO and the group of Diffusion-DPO [57] vs. Inversion-DPO. For each control group, we randomly select 500 pairs of images. In each trial, users are presented with two images and their corresponding prompts (text descriptions), one pair coming from SDXL and Inversion-DPO, and the other from Diffusion-DPO and Inversion-DPO. Users’ task is to choose the image that best matched the corresponding prompt. It is important to note that some of the generated images contain inappropriate content, so we remove a few images to ensure the experiment adhered to ethical guidelines and to prevent any potential negative impact.
The study invite 10 participants, with a gender ratio of 1:1, and ages ranging from 20 to 30. Participants come from diverse backgrounds, including computer science, design, and human-computer interaction (HCI).
User Preference Other Methods Inversion-DPO
SDXL vs Inversion-DPO 0.3347 0.6653
Diffusion-DPO vs Inversion-DPO 0.2733 0.7267
Table S4: Comparison of user preference between different image generation methods. In the comparison, Inversion-DPO consistently shows a higher preference from users, indicating its better alignment with human cognition and image content.
In the comparison between SDXL and Inversion-DPO, 66.53% of users select the image generated by Inversion-DPO, while only 33.47% chose the image generated by SDXL. In the comparison between Diffusion-DPO and Inversion-DPO, Inversion-DPO has a higher selection rate of 72.67%, while Diffusion-DPO has only 27.33%. Based on the results of these two control experiments, Inversion-DPO shows a clear user preference, whether compared with SDXL or Diffusion-DPO. Particularly in the comparison with Diffusion-DPO, the selection rate for Inversion-DPO was significantly higher, indicating that Inversion-DPO has a distinct advantage in terms of content alignment and consistency with human cognition when generating images.
Human Subjects Notification.Before the experiment began, we provided a notification to the participants to inform them about the collection and use of data.
Dear volunteers, thank you for your support of our research. We are studying an image generation algorithm that can generate high-quality images based on user-provided text prompts. All information about your participation in this study will be included in the study records. All information will be processed and stored according to the local privacy laws and policies. Your name will not appear in the final report. When referring to your data, only the individual number assigned to you will be mentioned.
The use of user data has been approved by the Institutional Review Board of the primary author’s affiliation.
S8 Additional Generation Examples
Figure S5: More high-quality and visually appealing generation examples.
Inversion-DPO demonstrates a highly competent capability in image generation, achieving visually impressive and high-quality results, as illustrated in Fig. S5. However, it is necessary to acknowledge that the model is not immune to failure cases. Typical failure modes include structural inconsistencies such as distorted legs, wings, or heads, as well as incorrect finger counts, as shown in the examples in Fig. S9.
S9 Additional Explanation of Efficiency
Figure S6: Relationship between training steps and Aesthetics Score improvements. The orange curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our implementation of Diffusion-DPO [57], while the dashed line corresponds to the officially released weights of Diffusion-DPO. All models are trained under identical configurations.
Figure S7: Relationship between training steps and CLIP Score improvements. The red curve shows the performance of our Inversion-DPO trained with different numbers of DDIM inversion steps. The gray curve represents our implementation of Diffusion-DPO [57], while the dashed line corresponds to the officially released weights of Diffusion-DPO. All models are trained under identical configurations.
In addition to the relationship between training steps and PickScore [26] presented in the main paper, we also report the correlations between training steps and Aesthetics Score [49] (Fig. S6), as well as between training steps and CLIP Score [44] (Fig. S7). Under identical training configurations, Inversion-DPO consistently outperforms Diffusion-DPO across all metrics, while also demonstrating faster convergence. Notably, even under constrained training conditions due to hardware limitations, Inversion-DPO still surpasses the officially released Diffusion-DPO in final performance.
S10 Visual Interpretation of the Underlying Assumptions
Figure S8: Visualization of data distributions in the inversion and sampling processes.
In our derivation, a key assumption based on DDIM inversion [52] is that the values of $\boldsymbol{x}_{t-1}^{w}$ are similar when obtained from the inversion of $x_{0}^{w}$ and from the forward sampling of $x_{T}^{w}$. To assess the validity of this assumption, we perform real-time visualization of the distributions along both the inversion and sampling trajectories. As illustrated in Fig. S8, the visual representations at corresponding time steps exhibit high similarity in both the forward and reverse processes, providing intuitive evidence that supports the reliability of this assumption.
S11 Detailed Derivation of InversionDPO
We present detailed derivation of the final loss of InversionDPO. We begin with the form as
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$ (S18)
Given the $\boldsymbol{x}_{1:T}^{w}$ by DDIM Inversion, we derive the following computation within the expectation. Specifically, $\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}$ is first decomposed into the differences in conditional probabilities at each time step:
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\left[\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})\right]$ (S19)
Since $p_{\theta}$ and $p_{\theta_{0}}$ share the same distribution of $x_{T}$ as standard Gaussian, $\log p_{\theta}(x_{T}^{w})-\log p_{\theta_{0}}(x_{T}^{w})$ cancel out. The conditional probabilities of $p_{\theta}$ and $p_{\theta_{0}}$ are Gaussian distributions as in equation 8 and becomes dirac-delta as $\sigma_{t}$ approaches 00.
Given the assumption of inversion in Sec. 3.2, the values of $\boldsymbol{x}_{t-1}^{w}$ are similar in the inversion from $x_{0}^{w}$ and in the sampling from $x_{T}^{w}$. Therefore, $\log p_{\theta}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ is only related to $\sigma_{t}^{2}$ and thus constant. At the same time, we have for $p_{\theta_{0}}$ its expecation is
$\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}+$ (S20)
$\left(\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}-\frac{\sqrt{\alpha_{t-1}({1-\alpha_{t}})}}{\sqrt{\alpha_{t}}}\right)\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)$ (S21)
Therefore, we have
$-\log p_{\theta_{0}}(\boldsymbol{x}_{t-1}^{w}|\boldsymbol{x}_{t}^{w})$ $=\frac{\|\boldsymbol{\mu}_{\theta_{0}}(x_{t}^{w})-\boldsymbol{x}_{t-1}^{w}\|}{2\sigma_{t}^{2}}^{2}+C$ (S22)
Since $x_{t-1}^{w}$ is equivalently generated by $p_{\theta}$ (i.e., $\mu_{\theta}$) based on the assumption above, we further assume $x_{t-1}^{w}$ subjects to $\mathcal{N}(\mu_{\theta}(x_{t}^{w}),\sigma_{t}^{2}I)$. By substituting it, we have the simplified form $\frac{1}{2\sigma_{t}^{2}}\|\boldsymbol{\mu}_{\theta}-\boldsymbol{\mu}_{\theta_{0}}\|^{2}$. Transforming the mean difference into the difference in epsilon prediction, and denoting the mean as $\boldsymbol{\mu}_{\theta}=\frac{1}{\sqrt{\alpha_{t}}}\boldsymbol{x}_{t}^{w}-\frac{\sqrt{1-\alpha_{t}}}{\sqrt{\alpha_{t}}}\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)$, we obtain:
$\|\boldsymbol{\mu}_{\theta}-\boldsymbol{\mu}_{\theta_{0}}\|^{2}=\frac{1-\alpha_{t}}{\alpha_{t}}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}$ (S23)
Summing the contributions across all time steps, we can arrive at:
$\log\frac{p_{\theta}(\boldsymbol{x}_{0:T}^{w})}{p_{\theta_{0}}(\boldsymbol{x}_{0:T}^{w})}=\sum_{t=1}^{T}\frac{1-\alpha_{t}}{2\sigma_{t}^{2}\alpha_{t}}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2},$ (S24)
We adopt the simple form as in DDPM [17] to eliminate the weighs and coefficients. Therefore, taking the winning and lossing sample all together, we reach the final loss as
$\mathcal{L}_{\text{Inversion-DPO}}(\theta)=-\mathbb{E}_{(\boldsymbol{x}_{0}^{w},\boldsymbol{x}_{0}^{l})}\log\sigma\bigg{(}\beta\mathbb{E}_{\begin{subarray}{c}\boldsymbol{x}_{1:T}^{w}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{w}|\boldsymbol{x}_{0}^{w})\\ \boldsymbol{x}_{1:T}^{l}\sim p_{\theta}(\boldsymbol{x}_{1:T}^{l}|\boldsymbol{x}_{0}^{l})\end{subarray}}$ (S25)
$\sum_{t=1}^{T}\Bigg{[}\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{w},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{w},t)\|^{2}-\|\epsilon_{\theta}(\boldsymbol{x}_{t}^{l},t)-\epsilon_{\theta_{0}}(\boldsymbol{x}_{t}^{l},t)\|^{2}\Bigg{]}\bigg{)}$
S12 Impact of the imprecise approximation
Model PickScore CLIP Aesthetic
Diffusion-DPO ($q$) 0.227 0.340 6.05
Inversion-DPO ($p_{\theta}$) 0.228 0.341 6.11
[…]
Table S5: Performance impact of approximating the reverse distribution $p_{\theta}$ with the forward $q$. In Inversion-DPO, the true reverse distribution $p_{\theta}$ is utilized, whereas Diffusion-DPO relies on its approximation $q$.
In Diffusion-DPO, due to the inaccessibility of the reverse trajectory, the posterior sampling $p_{\theta}(x_{1:T}|x_{0})$ is approximated by the noisy process $q(x_{1:T}|x_{0})$. We argue that this approximation introduces non-negligible training errors and conduct an ablation study to investigate this effect. While Inversion-DPO involves recovering the whole noise trajectory across multiple time steps for alignment, in the code implementation, the original Diffusion-DPO samples a single timestep from $q(x_{1:T}|x_{0})$ and augment with a value $T$ to obtain the final result. Such difference may also introduce performance discrepancy. For a fair comparison, we also evaluate Inversion-DPO by randomly sampling a single timestep from $p_{\theta}(x_{1:T}|x_{0})$. As presented in Tab. S5 in SDXL and SD1.5, training with DDIM Inversion with $p_{\theta}$ consistently yields better results than approximation based on the forward distribution $q$, highlighting the necessity and effectiveness of the core insight behind Inversion-DPO.
S13 A Brief Introduction of Stable Diffusion
Diffusion models, such as Denoising Diffusion Probabilistic Models (DDPMs) [17] and Denoising Diffusion Implicit Models (DDIMs) [52], provide a generative modeling framework that is based on the gradual corruption of data through a noise injection process followed by the recovery of the original data. These models leverage a Markov chain process to progressively add Gaussian noise to the data in a forward process, and then learn to reverse this process to recover the data from noisy samples.
In DDPM, the forward process is a Markov chain where data $x_{0}$ is progressively perturbed by Gaussian noise at each timestep, with the perturbation increasing over time. The transition between timesteps $x_{t-1}$ and $x_{t}$ is governed by a Gaussian distribution:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
where $\beta_{t}$ is the noise variance at timestep $t$, and the forward process gradually transforms the data $x_{0}$ into pure noise $x_{T}$. The reverse process, which is modeled during training, aims to recover the data by learning to reverse the noisy transformations:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)).$
The model is trained by optimizing the evidence lower bound (ELBO), which involves minimizing the difference between the forward and reverse process distributions. This training objective can be written as:
$\mathcal{L}_{\text{simple}}=\mathbb{E}_{x_{0},\epsilon,t}\left[\left\|\epsilon-\epsilon_{\theta}(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,t)\right\|^{2}\right],$
where $\epsilon_{\theta}$ is the model’s prediction of the noise added at each timestep, and $\epsilon$ is Gaussian noise. This loss function allows the model to learn to predict the noise at each timestep, effectively learning the reverse process.
To accelerate the sampling process, [52] introduces DDIMs, which extend DDPMs by generalizing the forward process to be non-Markovian. This modification allows for a more efficient sampling process that reduces the number of steps required to generate high-quality samples. In DDIMs, the generative process is deterministic, and the model can sample from fewer timesteps while maintaining high sample quality. The key insight behind DDIMs is that the reverse process in DDPM can be reparameterized using a deterministic trajectory, which simplifies the sampling process significantly.
S14 Limitation
Despite the effectiveness of our proposed Inversion-DPO framework, several limitations remain. First, the DDIM Inversion process relies on a strong assumption that the noise distribution in the forward (diffusion) and reverse (denoising) processes is consistent. Such assumption is valid only when the number of inversion steps is large enough. In practice, with a small inversion step, this assumption may potentially introduce inaccuracies in the reconstructed trajectories. Second, the training pipeline requires two rounds of inference: one from the pretrained model $p_{\theta_{0}}$ to obtain the DDIM-inverted noise trajectories, and another from the reference model $p_{\theta}$ to estimate the output distribution. This two inferences improve the convergence but increase training time and computational cost. Finally, the exact impact of the errors introduced by DDIM inversion remains unclear, and their influence on the overall optimization process warrants further investigation. More accurate Inversion methods may lead to more precision approximation. We present some failure cases in Fig. S9. Such failure may be a result of insufficient training or inadequate training data for aligment.
S15 Social Impact
Our work enhances preference optimization in diffusion models, offering improved controllability and sample quality. However, this also introduces potential risks. Automated preference signals, if not carefully designed, may inadvertently promote unsafe or biased content due to metric misalignment. Moreover, there exists the possibility of malicious misuse, where adversarial inputs could manipulate the optimization to favor harmful or misleading generations. We advocate for incorporating rigorous safety filters, human oversight, and transparent evaluation practices to mitigate such risks and ensure responsible deployment.
Figure S9: Additional results of failure examples.